OpenStudy (anonymous):

If f(x)=x^2/3, find f(-1) If f(x)=x^2/3, find f(-1) @Mathematics

jhonyy9 (jhonyy9):

1/3

OpenStudy (anonymous):

x^(2/3)= \[\sqrt[3]{x^2}\] so as (-1)^2=1 and \[\sqrt{1}\]=1 (-1)^(2/3)=1

OpenStudy (anonymous):

nevermind im wrong http://www.wolframalpha.com/input/?i=%28-1%29%5E%282%2F3%29

jhonyy9 (jhonyy9):

2/3 want to be exponent ???

jhonyy9 (jhonyy9):

indifferent this will be equal 1

OpenStudy (anonymous):

yes, you were correct, 1 is the correct answer

OpenStudy (anonymous):

oooh lol if not then its super easy :)

jhonyy9 (jhonyy9):

but in future in case of this style of exercise will be right,correct if you will us paratheses